# Towards bandit-based prompt-tuning for in-the-wild foundation agents

### v2

@@ 4.1 Baselines
Standard PDT [Xu et al., 2022] without prompt-tuning: This baselines reveals the possible performance gains due to prompt-tuning at inference time.
ZO-RankSGD-based prompt-tuning [Hu et al., 2023]: Closely related to our bandit-based method, this approach proposes prompt-tuning for PDT by employing ZO-RankSGD [Tang et al., 2023] to estimate the gradient of the prompt with respect to online task return $G$. The method samples and initial prompt $\rho_{0}\sim\mathcal{P}_{i}$, and, at each rounds $k$, estimates the gradient ^$\hat{\nabla}_{\rho}G$ based on the ranking between $m$ perturbed versions of $\rho_{k}$. The perturbed versions of the prompt are obtained as $\rho_{k}^{\prime}=\rho_{k}+\epsilon\mathcal{N}(0,I_{d})$, where $\epsilon$ is the noise scale, $I_{d}$ is the $d\times d$ identity matrix, and $d=|\rho|$ is the length of the prompt. The prompt is then updated according to $\rho_{k+1}\leftarrow\rho_{k}+\eta\hat{\nabla}_{\rho}G$, where $\eta$ is the learning rate that we anneal from $1$ to $0.1$ over the $K$ rounds. Crucially, at each round $k$, all of the $m$ prompt-perturbations must be evaluated with an online rollout of the PDT, meaning the sample complexity of this method is $m$ times larger that of our method.
@@ 4.2 Results & Analysis
Does bandit-based prompt-tuning improve a frozen PDT backbone? We perform prompt tuning on the pre-trained PDT $\theta^{*}$ using 250 online rollouts on training tasks with radius $r=2.9$. We run this experiment with $J\in\{1,2,4\}$, i.e., with increasingly many segments and, conversely, tokens for task identification in the prompt. We run our bandit-based prompt-tuning method with UCB [Li et al., 2010] and $\epsilon$-greedy exploration strategies.
@@ 5 Conclusion & Future Work
We introduce a bandit-based prompt-tuning approach for multi-task foundation agents which efficiently navigates the prompt space to identify high-performing prompts for downstream tasks, avoiding exhaustive random sampling. Preliminary results in a 2D proof-of-concept environment suggest that our method enhances a pre-trained PDT to optimal performance where uniform prompt sampling fails. We believe this warrants future research and are exploring the scalability of the proposed method to more complex and challenging environments and aim to present additional results at the workshop.
A potential limitation of our approach is the use of independent reward models for the $J$ segments in the stochastic trajectory prompt, which neglects correlations among segments and leads to a high-variance estimator. An important future direction is to develop a more sophisticated bandit architecture that disentangles segment contributions to overall performance.

### v3

@@ 4.1 Baselines
Standard PDT [19] without prompt-tuning: This baselines reveals the possible performance gains due to prompt-tuning at inference time.
ZO-RankSGD-based prompt-tuning [6]: Closely related to our bandit-based method, this approach proposes prompt-tuning for PDT by employing ZO-RankSGD [17] to estimate the gradient of the prompt with respect to online task return $G$. The method samples and initial prompt $\rho_{0}\sim\mathcal{P}_{i}$, and, at each rounds $k$, estimates the gradient ^$\hat{\nabla}_{\rho}G$ based on the ranking between $m$ perturbed versions of $\rho_{k}$. The perturbed versions of the prompt are obtained as $\rho_{k}^{\prime}=\rho_{k}+\epsilon\mathcal{N}(0,I_{d})$, where $\epsilon$ is the noise scale, $I_{d}$ is the $d\times d$ identity matrix, and $d=|\rho|$ is the length of the prompt. The prompt is then updated according to $\rho_{k+1}\leftarrow\rho_{k}+\eta\hat{\nabla}_{\rho}G$, where $\eta$ is the learning rate that we anneal from $1$ to $0.1$ over the $K$ rounds. Crucially, at each round $k$, all of the $m$ prompt-perturbations must be evaluated with an online rollout of the PDT, meaning the sample complexity of this method is $m$ times larger than that of our method.
@@ 4.2 Results & Analysis
Does bandit-based prompt-tuning improve a frozen PDT backbone? We perform prompt tuning on the pre-trained PDT $\theta^{*}$ using 250 online rollouts on training tasks with radius $r=2.9$. We run this experiment with $J\in\{1,2,4\}$, i.e., with increasingly many segments and, conversely, tokens for task identification in the prompt. We run our bandit-based prompt-tuning method with UCB [9] and $\epsilon$-greedy exploration strategies.
@@ 5 Conclusion & Future Work
We introduce a bandit-based prompt-tuning approach that enables few-shot generalization at inference time for multi-task decision transformers. By efficiently exploring the prompt space, our approach identifies high-performing prompts for downstream tasks, avoiding exhaustive random sampling or full-model finetuning.
Our preliminary results show that the proposed contextual bandit architecture efficiently boosts a pre-trained PDT to optimal levels of performance, after only a small number of online rollouts, while uniform sampling and prior prompt-tuning approaches underperform.
We believe these findings warrant future research and are exploring the scalability of the proposed method to more complex and challenging environments and aim to present additional results at the workshop.